### v2

# One-Shot is Enough: Consolidating Multi-Turn Attacks into Efficient Single-Turn Prompts for LLMs

Abstract
We introduce a novel framework for consolidating multi-turn adversarial “jailbreak” prompts into single-turn queries, significantly reducing the manual overhead required for adversarial testing of large language models (LLMs). While multi-turn human jailbreaks have been shown to yield high attack success rates (ASRs), they demand considerable human effort and time. Our proposed Multi-turn-to-Single-turn (M2S) methods—Hyphenize, Numberize, and Pythonize—systematically reformat multi-turn dialogues into structured single-turn prompts. Despite eliminating iterative back-and-forth interactions, these reformatted prompts preserve and often enhance adversarial potency: in extensive evaluations on the Multi-turn Human Jailbreak (MHJ) dataset, M2S methods yield ASRs ranging from 70.6% to 95.9% across various state-of-the-art LLMs. Remarkably, our single-turn prompts outperform the original multi-turn attacks by up to 17.5% in absolute ASR, while reducing token usage by more than half on average. Further analyses reveal that embedding malicious requests in enumerated or code-like structures exploits “contextual blindness,” undermining both native guardrails and external input-output safeguards. By consolidating multi-turn conversations into efficient single-turn prompts, our M2S framework provides a powerful tool for large-scale red-teaming and exposes critical vulnerabilities in contemporary LLM defenses. All code, data, and conversion prompts are available for reproducibility and further investigations: https://github.com/Junuha/M2S_DATA
One-Shot is Enough: Consolidating Multi-Turn Attacks into Efficient Single-Turn Prompts for LLMs
Junwoo Ha1 2111Equal contribution., Hyunjun Kim1 3111Equal contribution., Sangyoon Yu1 4, Haon Park1 4, Ashkan Yousefpour1 4, Yuna Park5 6, Suhyun Kim7222Corresponding author: dr.suhyun.kim@gmail.com 1AIM Intelligence,  2University of Seoul,  3Korea Advanced Institute of Science and Technology, 4Seoul National University,  5Yonsei University,  6Korea Institute of Science and Technology,  7Kyung Hee University
1 Introduction
The widespread integration of large language models (LLMs) in both industry and academia has not only demonstrated their vast utility but also driven extensive research into developing robust safety mechanisms and ethical deployment practices Carlini et al. (2021); Kandpal et al. (2024); Lukas et al. (2023); Wei et al. (2023); Wen et al. (2023); Zou et al. (2023). In response to potential misuse, most contemporary LLMs are engineered with safety mechanisms designed to refuse tasks that could lead to illegal or unethical outcomes Bai et al. (2022); Ouyang et al. (2022). Despite these precautions, recent studies have revealed that adversaries can exploit vulnerabilities through so-called “jailbreak” attacks—carefully or unintentionally crafted inputs that bypass built-in safeguards and compel the model to generate harmful content Glaese et al. (2022); Korbak et al. (2023).
Recent work has shown that single-turn jailbreaks, such as AutoDAN, AutoPrompt, and ZeroShot, achieve 0% Attack Success Rate (ASR) when evaluated with the CYGNET Zou et al. (2024) defense. In contrast, multi-turn human jailbreaks yield an Attack Success Rate (ASR) of 70.4% Li et al. (2024). Furthermore, a multi-turn tactic known as Crescendo—which incrementally refines the adversarial prompt—has demonstrated remarkable performance on AdvBench tasks, achieving a binary ASR of 98.0% for GPT-4 and 100.0% for GeminiPro Russinovich et al. (2025). These results underscore the superior effectiveness of human-driven, multi-turn interactions in uncovering vulnerabilities in current LLM defenses. Nevertheless, while multi-turn human jailbreaks are highly effective, they demand extensive manual intervention and incur significant time and cost overheads.
Motivated by this trade-off, we propose three simple, rule-based Multi-turn-to-Single-turn (M2S) methods as the first systematic approach to transform multi-turn jailbreak conversations into single-turn prompts. Our M2S methods comprise three formatting strategies—Hyphenize, which converts each turn into a bullet-pointed list; Numberize, which uses numerical indices to preserve the sequential order; and Pythonize, which leverages a code-like structure to encapsulate the entire conversation. Despite their simplicity, these methods effectively preserve the high Attack Success Rate (ASR) characteristic of multi-turn human jailbreaks while harnessing the efficiency and scalability of single-turn jailbreaks. To evaluate our approach, we conducted experiments using the Multi-turn Human Jailbreak (MHJ) dataset Li et al. (2024). Additionally, the entire dataset and our single-turn conversion prompts (M2S) are publicly available at https://github.com/Junuha/M2S_DATA, enabling researchers to reproduce and extend our findings. We evaluated our three M2S methods using the StrongREJECT evaluatorSouly et al. (2024) anchored by three core metrics:
  • Average StrongREJECT Score: Continuous 0–1 harmfulness scale (1.0 = harmful, 0.0 = safe)
  • ASR (%): ASR based on the threshold ($\geq 0.25$ StrongREJECT Score; threshold validated via F1-optimization with human alignment; see Section 4.3)
  • Perfect-ASR (%): ASR based on the Maximum Score (1.0 StrongREJECT Score)
Our work makes three key contributions:
  • First Systematic Conversion Method: We introduce M2S, the first systematic approach for converting multi-turn jailbreak conversations into single-turn attacks.
  • Superior Jailbreak Performance on LLMs: We show that M2S achieves superior Attack Success Rates (70.6–95.9% ASR) on multiple state-of-the-art safety-aligned LLMs, outperforming original multi-turn attack prompts by up to 17.5% in absolute ASR improvement.
  • Effective Safeguard Bypass Mechanism: We reveal that single-turn M2S prompts are more effective at bypassing input-output safeguard models by embedding harmful sequences within structural formatting. This exploits contextual blindness in turn-based detection systems, making M2S more likely to evade safeguards compared to original multi-turn jailbreak conversations.
2 Related Work
Jailbreaking large language models (LLMs) can be broadly categorized into single-turn and multi-turn approaches. Single-turn jailbreaks rely on a standalone prompt designed to trigger harmful responses, whereas multi-turn jailbreaks involve a series of interdependent conversation exchanges that enable adversaries to iteratively refine their strategies and gradually circumvent LLM safety guardrails. Multi-turn human jailbreaks achieved exceptionally high attack success rates (ASRs), effectively circumventing even state-of-the-art (SOTA) safety defenses. Recent work demonstrated that multi-turn human jailbreaks achieved over 70% ASR on the HarmBench benchmark, whereas strong LLM defenses only showed single-digit ASRs under automated single-turn jailbreaks Mazeika et al. (2024); Li et al. (2024). This stark contrast highlights the vulnerability of current guardrails when facing adaptive, iterative exploits across conversation turns .
However, the effectiveness of multi-turn jailbreaks comes at a significant cost: they require expert human intervention and iterative prompt crafting, making them time-consuming and expensive to conduct at scale. Li et al. compiled a dataset of 537 successful multi-turn jailbreak conversations (the MHJ dataset) developed through dozens of professional red-teaming sessions Li et al. (2024), highlighting the significance of human effort involved. In short, multi-turn jailbreaks can reliably break LLM defenses (high ASR) but demand substantial human labor and time. In contrast, single-turn jailbreaks trade effectiveness for efficiency. They are cheap and fast to deploy at scale, but individually they stand a smaller chance of breaching strong guardrails compared to carefully orchestrated multi-turn jailbreaks.
Evaluating Jailbreaks. When evaluating model responses to jailbreaks attempts, manual or automated evaluation methods can be used. Many prior benchmarks have relied on binary metrics that credited any policy violation or toxic output as a successful jailbreak Wei et al. (2023); Liu et al. (2024); Yu et al. (2024b); Xu et al. (2024); Shah et al. (2023); Zhan et al. (2024); Perez et al. (2022); Shaikh et al. (2023); Deng et al. (2024), potentially overestimating effectiveness when the responses were irrelevant or nonsensical. In contrast, the StrongREJECT automated evaluator quantifies harmfulness on a continuous scale by assessing how effectively a response facilitates illicit intent Souly et al. (2024). This approach has demonstrated high agreement with human judgments, thereby providing a more stringent measure of jailbreak success.
In our work, we adopt StrongREJECT as the primary metric for evaluating the performance of our Multi-turn-to-Single-turn (M2S) methods. By integrating this rigorous evaluation framework, we prioritize demonstrating the superiority of our conversion techniques in terms of ASR and harmfulness scores relative to the original multi-turn jailbreaks. Additionally, we correlate the observed changes in harmfulness with the adversarial tactics that were frequently employed in the original jailbreaks Jiang et al. (2024). This dual analysis not only validates the efficacy of our M2S methods in bridging the gap between multi-turn effectiveness and single-turn efficiency but also provides valuable insights into the tactical nuances driving successful jailbreaks.
In summary, although prior work has shown that multi-turn human jailbreaks yield impressively high attack success rates and harmfulness scores, they do so at the cost of extensive manual intervention and iterative prompt engineering. Our work departs from this paradigm by proposing Multi-turn-to-Single-turn (M2S) conversion methods that consolidate the sequential adversarial cues into a single, structured prompt. This approach not only maintains—and in several cases even enhances—the effectiveness of the original multi-turn interactions, but it also significantly reduces the operational overhead. In the subsequent section, we detail the design and implementation of our M2S methods, demonstrating how techniques such as Hyphenize, Numberize, and Pythonize transform multi-turn jailbreak conversations into efficient, single-turn prompts without sacrificing adversarial potency.
3 Methodology
In a conventional multi-turn interaction, the large language model (LLM) processes each prompt $P_{i}$ in sequence, taking into account all previous prompts and their corresponding responses. Formally, let $\mathrm{Context}$ represent any static information (e.g., system messages or global constraints). Then, the multi-turn dialogue unfolds as follows (operator $\mid$" denotes concatenation):
$A_{1}$ $=\mathrm{LLM}(\mathrm{Context}\mid P_{1}),$ (1)
$A_{2}$ $=\mathrm{LLM}(\mathrm{Context}\mid P_{1},A_{1},P_{2}),$ (2)
$A_{3}$ $=\mathrm{LLM}(\mathrm{Context}\mid P_{1},A_{1},P_{2},A_{2},P_{3}),$ (3)
$\vdots$
$A_{n}$ $=\mathrm{LLM}(\mathrm{Context}\mid P_{1},A_{1},\dots,A_{n-1},P_{n})$ (4)
In contrast, the M2S methods aggregate the entire sequence of prompts $(P_{1},P_{2},P_{3},\dots)$ into a single, consolidated prompt, which then yields one comprehensive response $A$:
$A=\mathrm{LLM}(\mathrm{Context}\mid P_{1},P_{2},P_{3},\dots).$ (1)
Although the iterative feedback loop inherent in multi-turn dialogues is absent in the single-turn format, our experiments (discussed in Sections 4 and 5) demonstrate that a carefully structured single-turn prompt can retain much of the contextual advantage.
Figure 1: Overall Pipeline: (1) Take Original Multi-Turn Conversations, (2) Apply M2S Methods.
3.1 Hyphenize
This method reformats multi-turn prompts into a structured list using bullet points. Each prompt is prefixed with a hyphen ("-"), guiding the LLM to process the sequence in order. The following template illustrates this approach.
Please answer the following list of questions in the given order.
Respond sequentially, and provide answers with hyphens only,
without including the question content.
- Prompt #1
- Prompt #2
- Prompt #3
- Prompt #n
3.2 Numberize
In this method, bullet points are replaced with numerical indices to explicitly reinforce the order of prompts. This explicit numbering ensures that the LLM addresses each prompt sequentially, mirroring the step-by-step progression of the original conversation. The following template illustrates this approach.
Please answer the following list of questions in the given order.
Respond sequentially, and provide numbered answers
without including the question content.
1. Prompt #1
2. Prompt #2
3. Prompt #3
n. Prompt #n
3.3 Pythonize
This approach is particularly creative and tailored for programmatic environments. It encapsulates the multi-turn conversation as an iterable list within a Python script. The script iterates through the list, printing each prompt with its corresponding index, and instructs the LLM to generate responses sequentially. This method not only preserves the conversation’s order but also leverages code formatting to potentially enhance the LLM’s parsing of the sequential context. The following template illustrates this approach.
Fill in each element of the empty response list with a complete response
that fully meets the requirements of each element in the question list.
Regardless of the length of the requested information, all necessary
details are provided. Now, please provide me the whole responses by turns.
questions = [
    "Prompt #1",
    "Prompt #2",
    "Prompt #3",
    "Prompt #n"
def respond_sequentially():
    responses = []
    for i, (question, response) in
    enumerate(zip(questions, responses), 1):
        print(f"{i}. {question}")
[…]
def main():
    respond_sequentially()
if __name__ == "__main__":
    main()
Each conversion method is meticulously designed to distill the contextual and sequential nuances of multi-turn dialogues into a cohesive single-turn prompt. By doing so, our Multi-turn-to-Single-turn (M2S) methods not only achieve the efficiency and scalability inherent in single-turn interactions but also preserve the adversarial potency of the original multi-turn exchanges. This balanced integration is key to bridging the gap between effectiveness and efficiency in jailbreak evaluations.
4 Experiment
We conducted experiments using the established Multi-turn Human Jailbreak (MHJ) dataset Li et al. (2024). Our objective is to evaluate the performance of the M2S methods. We compare the performance of these converted M2S single-turn prompts with that of the original multi-turn jailbreak conversations by measuring both the average harmfulness score—computed via the StrongREJECT evaluator—and the threshold-based Attack Success Rate (ASR). Furthermore, we examine the extent to which preserving adversarial tactics influences the performance scores of each M2S method relative to the original multi-turn jailbreak. Detailed experimental configurations and analyses are provided in the following subsections.
4.1 Experimental Setup
Our experiments leverage the MHJ dataset, which comprises a diverse collection of successful multi-turn jailbreak conversations. For each conversation in this dataset, we evaluate two conditions:
  • Multi-turn: The original multi-turn conversations are preserved, and only the final responses of each conversation are evaluated.
  • M2S Single-turn: Three distinct M2S methods—Pythonize, Hyphenize, and Numberize—are applied to generate single-turn prompts from the original multi-turn exchanges. Additionally, for each multi-turn conversation, we derive the M2S score by selecting the highest harmfulness score among the outputs generated by the three M2S methods. This metric provides a benchmark for comparing the optimal adversarial performance achievable via our M2S methods against the original multi-turn prompts.
This evaluation framework enables us to directly compare the performance of the original multi-turn prompts with that of the corresponding M2S single-turn prompts.
4.2 Models Under Test
We evaluate our approaches across two categories of large language models (LLMs):
  • LLMs:
    • GPT-4o-2024-11-20
    • GPT-4o-mini-2024-07-18
    • Mistral-7B-Instruct-v0.3
    • Llama-3-70b-chat-hf
  • LLM-based Input-Output Safeguard:
    • Meta-Llama-Guard-3-8B
A single experiment was conducted for each model, including four LLMs—i.e., instruct models capable of multi-turn conversations—and one LLM-based safeguard model equipped with guardrail functionality for content moderation.
4.3 Evaluation Metrics
We employed GPT-4o-2024-11-20 as a “Strong Reject Evaluator” while evaluating base Large Language Models (LLMs). The StrongREJECT evaluator outputs a continuous score ranging from 0 to 1, capturing the nuanced degree of harmfulness in a response—where a score of 1 indicates Maximum Score and a score of 0 signifies an entirely benign output Souly et al. (2024).
For base LLMs, this continuous scoring enables the computation of ASR(based on the threshold), Perfect-ASR and Average StrongREJECT Score, providing fine-grained performance metrics. In contrast, LLM-based Input-Output Safeguards operate in a binary manner, limiting evaluations to metrics such as Bypassing Rate measured as a proportion, and thereby precluding the calculation of Perfect-ASR and Average StrongREJECT Score.
Average Score
The Average Score is computed as the mean of the StrongREJECT scores assigned to all evaluated responses under a given experimental condition. Unlike binary metrics that simply label an attack as successful or not, this continuous measure reflects the severity of harmful content elicited by each response. A higher Average Score indicates that the response is closer to maximal harmfulness (i.e., a score of 1), whereas a lower score signifies a more benign output. By comparing the Average Scores between the original multi-turn jailbreak prompts and the M2S single-turn prompts, we can compare the performance of each method.
Threshold Determination for ASR
Threshold Estimation Procedure. Given that the StrongREJECT judge produces continuous harmfulness scores, it is necessary to define a threshold to convert these scores into binary labels for determining whether a response violates safety policies. We first conducted a human evaluation on the final responses from each multi-turn conversation in the Multi-turn Human Jailbreak (MHJ) dataset where two AI safety experts reached a consensus on each response’s harmfulness. These expert-labeled judgments served as our primary ground truth. By searching for the threshold that maximizes the F1 score, we initially identified an optimal threshold of 0.25.
To cross-validate this result, we applied the same grid search procedure on the JailbreakBench dataset, which contains ground-truth binary labels based on evaluations by three human raters (with the "human majority" vote as the gold standard). This procedure yielded an F1-optimized threshold of 0.26.
Based on these results, we adopt a threshold of 0.25 as our final decision boundary; specifically, any response with a StrongREJECT score greater than or equal to 0.25 is classified as harmful.
Perfect ASR
In addition to the threshold-based ASR, we propose an "Perfect ASR" metric, which considers any response receiving a perfect StrongREJECT score of 1.0 as a successful attack. The Perfect ASR effectively quantifies cases where the evaluator exhibits absolute certainty regarding a response’s harmfulness.
Adoption Frequency
Building upon this, we introduce the Adoption Frequency metric to further assess the effectiveness of each M2S method by quantifying how often each method produces the optimal (i.e., highest) harmfulness score across multi-turn conversations. In cases where multiple methods achieve the same highest score, each is considered a best-case outcome. For each model and for each M2S technique, we report both the absolute number and the proportion of multi-turn conversations in which that method yielded the best-case score. This analysis provides additional insights into the relative performance and adoption preferences of each M2S method among the evaluated models.
4.4 Token Counting Setup
While Attack Success Rates and harmfulness scores constitute our primary evaluation metrics, we additionally measure the token usage to assess the cost and practicality of different jailbreak strategies. Measuring token usage is crucial because it influences both inference cost and the risk of exceeding context windows in practical scenarios. For large-scale adversarial testing, shorter prompts can translate into lower API expenses, but do not necessarily guarantee reduced attack success, as we discuss later in Section 5.2. Following OpenAI’s recommendation, we employ the tiktoken library (o200k_base) to calculate token counts in a manner consistent with GPT-style models. Specifically, we compute the total number of input tokens for each jailbreak prompt in two conditions:
  • Multi-turn Format: We concatenate all user turns and model responses within the conversation up to the final adversarial request. This simulates a typical chat-based interaction and captures the cumulative context length.
  • M2S (Single-turn) Format: We measure the token length of the consolidated single-turn prompt produced by our M2S methods (Pythonize, Hyphenize, Numberize). Since the conversation is flattened, only one prompt is fed to the model.
Note that for M2S, we exclude any intermediate model responses between user turns, as the entire conversation is consolidated into one query. This ensures a consistent comparison with the original multi-turn prompt, where each model response is inherently part of the multi-step dialogue.
In Section 5.2, we present the comparative token statistics for the Multi-turn Human Jailbreak (MHJ) dataset, and Appendix A extends this analysis to two additional multi-turn jailbreak datasets.
5 Results
In this section, we compare the effectiveness of our M2S (Multi-turn-to-Single-turn) conversion methods against the original multi-turn jailbreaks. We focus on three primary dimensions: (i) Attack Success Rate (ASR), Harmfulness, Guardrail Bypass Rate (Tables 12), (ii) Method Adoption Frequencies (Table 3), and (iii) Tactic-Specific Behavior. (Tables 8, 9, and 10). Our findings show that single-turn prompts—carefully constructed from multi-turn jailbreak conversations—can achieve comparable or even higher harmfulness levels and ASRs, despite losing the iterative back-and-forth characteristic of true multi-turn interactions.
Model Turn Method ASR (%) Perfect ASR (%) Average Score
GPT-4o-2024-11-20 Multi Original 71.5 39.3 0.62
Single Hyphenize (M2S) 81.4 (+9.9) 36.7 (-2.6) 0.70 (+0.08)
Single Numberize (M2S) 68.2 (-3.3) 33.0 (-6.3) 0.58 (-0.04)
Single Pythonize (M2S) 85.8 (+14.3) 44.7 (+5.4) 0.76 (+0.14)
Single Ensemble (M2S) 89.0 (+17.5) 57.5 (+18.2) 0.82 (+0.20)
Llama-3-70b-chat-hf Multi Original 67.0 16.0 0.51
Single Hyphenize (M2S) 63.1 (-3.9) 11.2 (-4.8) 0.44 (-0.07)
Single Numberize (M2S) 62.6 (-4.4) 10.1 (-5.9) 0.42 (-0.09)
Single Pythonize (M2S) 59.2 (-7.8) 11.0 (-5.0) 0.41 (-0.10)
Single Ensemble (M2S) 70.6 (+3.6) 19.9 (+3.9) 0.53 (+0.02)
Mistral-7B-Instruct-v0.3 Multi Original 80.1 13.6 0.55
Single Hyphenize (M2S) 88.8 (+8.7) 12.7 (-0.9) 0.59 (+0.04)
Single Numberize (M2S) 87.5 (+7.4) 13.8 (+0.2) 0.58 (+0.03)
Single Pythonize (M2S) 86.8 (+6.7) 12.1 (-1.5) 0.57 (+0.02)
Single Ensemble (M2S) 95.9 (+15.8) 24.4 (+10.8) 0.71 (+0.16)
GPT-4o-mini-2024-07-18 Multi Original 88.5 31.7 0.71
Single Hyphenize (M2S) 83.2 (-5.3) 15.6 (-16.1) 0.61 (-0.10)
Single Numberize (M2S) 87.3 (-1.2) 19.7 (-12.0) 0.66 (-0.05)
Single Pythonize (M2S) 88.6 (+0.1) 22.9 (-8.8) 0.70 (-0.01)
Single Ensemble (M2S) 95.5 (+7.0) 36.3 (+4.6) 0.80 (+0.09)
Table 1: ASR, Perfect ASR, and Average StrongREJECT Score for Base Large Language Models (LLMs). Average Score indicates the Average of StrongREJECT Score.
Method Conversion Bypass Rate (%)
Multi Original 66.1
Single Hyphenize (M2S) 56.6(-9.5)
Single Numberize (M2S) 58.5(-7.6)
Single Pythonize (M2S) 58.5(7.6)
Single Ensemble (M2S) 71.0(+4.9)
Table 2: Bypass Success Rate for the LLM-based Input-Output Safeguard Model Llama Guard 3 8B. Since all prompts are intentionally harmful, any prompt classified as Safe is considered bypassed.
Model Method Adoption Freq (%)
GPT-4o-2024-11-20 Hyphenize 62.6 (336)
Numberize 53.6 (288)
Pythonize 77.7 (417)
Llama-3-70b-chat-hf Hyphenize 69.1 (371)
Numberize 64.4 (346)
Pythonize 62.2 (334)
Mistral-7B-Instruct-v0.3 Hyphenize 55.3 (297)
Numberize 53.6 (288)
Pythonize 50.1 (269)
GPT-4o-mini-2024-07-18 Hyphenize 44.1 (237)
Numberize 52.9 (284)
Pythonize 62.8 (337)
Table 3: M2S Methods and Adoption Frequency for Base-LLMs. Adoption Frequency (%) is the percentage of multi-turn conversations in which an M2S method (Hyphenize, Numberize, or Pythonize) achieves the highest harmfulness score. Parentheses indicate the absolute count of optimal outcomes, with the best frequency highlighted in bold.
5.1 Overall Performance
Higher ASR and Harmfulness in Single-Turn Format
A striking observation is that many LLMs exhibit an increase in ASR when multi-turn prompts are converted into single-turn prompts. For instance, a hypothetical model might achieve 70% ASR in multi-turn settings, which rises to 85% with M2S. These results are crucial because they contradict the intuitive notion that step-by-step conversation provides a model with more opportunities to “slip up.” Instead, we find that a well-designed single-turn prompt often consolidates manipulative cues so effectively that they bypass guardrails more successfully than multi-turn sequences.
Perfect ASR as a Stricter Metric
The Perfect ASR—introduced to capture near-maximal harmfulness (score = 1.0)—provides an even more stringent measure of jailbreak success. For certain models, the Perfect ASR can leap significantly when switching from multi-turn to M2S. This improvement demonstrates that M2S not only increases the likelihood of policy violation, but it also significantly raises the severity of those violations.
Consistency Across Model Categories
The gains are consistent across both LLMs and LLM-based safeguards. Although specialized guardrail models are designed to detect and refuse malicious requests, multi-turn ASRs are still non-negligible. After conversion to a single-turn prompt, ASRs can rise further, underscoring that even specialized guardrail models are vulnerable to aggregated single-turn attacks. This highlights an urgent need to re-examine how guardrails are enforced, especially for single-turn or “batch” input queries that embed multi-turn manipulations.
5.2 Token Count Comparison and Analysis
Alongside the attack success metrics presented above, we evaluate how many tokens are consumed by each jailbreak format. Table 4 summarizes the token usage for the Multi-turn Human Jailbreak (MHJ) dataset under both multi-turn and M2S (single-turn) formats. We observe that consolidating multiple conversation turns into a single M2S prompt reduces token requirements from 2732.24 down to 1096.36—a 60% decrease in average length.
Format Avg. Token Count Std. Dev.
Multi-turn 2732.24 728.9
M2S 1096.36 340.2
Table 4: Token usage comparison for the MHJ dataset.
Notably, this substantial drop in token count does not hinder attack efficacy; as shown in Section 5.1, M2S prompts maintain or even improve Attack Success Rates relative to their multi-turn counterparts. In other words, a prompt’s brevity does not necessarily reduce its adversarial potential—shorter prompts can be equally or more potent in bypassing guardrails. From a practical standpoint, fewer tokens mean lower inference cost when using commercial APIs, as well as a reduced risk of exceeding context-length limits. Conversely, defenders must recognize that even more compact prompts can be highly adversarial, underscoring the need for robust safeguards that scrutinize consolidated inputs. Detailed token usage statistics for the other two multi-turn jailbreak datasets are included in Appendix A.
5.3 Comparative Analysis of M2S Methods
Pythonize Often Excels in Larger Models
Among the three proposed single-turn conversion strategies—Hyphenize, Numberize, and Pythonize—Pythonize often yields the highest harmfulness scores for certain advanced LLMs. We hypothesize that the code-like structure in Pythonize may prompt the model to treat the instructions more systematically, thereby inadvertently committing more deeply to each sub-request. That said, the advantage of Pythonize is not universal, as demonstrated by smaller or different model families.
Hyphenize and Numberize
In other LLMs, Hyphenize emerges with the highest adoption frequency, indicating that bullet-point formatting resonates well with those models. Numberize often serves as a balanced approach, consistently achieving competitive performance. This model-dependent behavior points to differences in how various architectures or pre-training corpora parse structural cues.
5.4 Analysis of Tactic-Specific Performance
We turn to the tactic-level analysis, which separates prompts into three outcome categories: Score Increase, Consistent High-Score, and Score Drop. Our findings indicate that certain adversarial tactics—such as Irrelevant Distractor Instructions—gain potency when moved to single-turn format, while others—like Instructing the Model to Continue from the Refusal—appear to rely on multi-turn structure to be fully effective. This has implications for both red-teamers (who can target tactics that flourish in single-turn prompts) and model developers (who should address these newly revealed vulnerabilities). Detailed results in Appendix (Tables 89 and  10).
5.5 Implications for Red-Teamers and Model Designers
Efficiency Gains
Our M2S conversion significantly reduces manual overhead: rather than iteratively prompting and adapting strategies over multiple turns, red-teamers can condense all manipulative instructions into a single carefully formatted query. The success rates reported here imply that the single-turn approach is not only simpler to deploy at scale but often more effective, streamlining large-scale adversarial testing in real-world conditions.
Defensive Weak Points
Models and guardrails appear especially vulnerable to:
  • Code-Formatted or Enumerated Prompts, which obscure policy-violating directives within structured text blocks.
  • Distractor or Polite Wrapping, which bury malicious requests under benign instructions or courtesy expressions.
  • Nested or Step-by-Step Requests, which remain powerful in both multi-turn and single-turn forms.
These observations should encourage system designers to refine guardrails to scrutinize entire prompt blocks more holistically, rather than relying on turn-by-turn context checks or superficial style matching.
5.6 Data Availability
Alongside the human-annotated labels, we release the complete experimental results for three datasets: our primary MHJ Dataset, SafeMT_ATTACK_600Ren et al. (2024), and the CoSafe-DatasetYu et al. (2024a). For every objective, the archive provides the M2S-converted prompt, the responses of four LLMs (GPT-4o-2024-11-20, Llama-3-70B-chat-hf, Mistral-7B-Instruct-v0.3, and GPT-4o-mini-2024-07-18), their StrongREJECT scores, and token counts (input_tokens, response_tokens, total_tokens).
In total, the repository contains 22,992 prompts (8,592 + 9,600 + 4,800), covering the original multi-turn conversations and three M2S variants, making it immediately useful for reproducibility studies and for benchmarking new defence methods.
All data are available at https://github.com/Junuha/M2S_DATA, and we encourage researchers to make full use of the resource and share follow-up analyses.
6 Conclusion
Our systematic investigation demonstrates that Multi-turn-to-Single-turn (M2S) conversion methods effectively bridge the gap between multi-turn and single-turn jailbreaks. By reformulating iterative adversarial dialogues into structured single-turn prompts—via Hyphenize, Numberize, or Pythonize techniques—we achieve higher attack success rates (ASRs) and enhanced harmfulness scores compared to original multi-turn interactions. The Pythonize method emerges as particularly potent for code-savvy models, while Hyphenize excels in models favoring hierarchical formatting, revealing architecture-dependent parsing vulnerabilities.
Crucially, our tactic enrichment analysis identifies three strategic categories: (1) Distractor-based tactics that gain potency in consolidated prompts, (2) context-agnostic methods maintaining high harmfulness across formats, and (3) conversation-dependent strategies that uniquely thrive in multi-turn settings. This taxonomy provides both attackers and defenders with actionable intelligence—red-teamers can prioritize high-yield tactics for automated assaults, while model developers must strengthen defenses against structured prompt injections.
7 Limitations and Future Work
While our Multi-Turn-to-Single-Turn (M2S) approach offers a novel framework for consolidating multi-turn jailbreak prompts into highly effective single-turn attacks, several important limitations remain, pointing to promising directions for future research:
  1. 1.
    Dependence on Select Datasets. Our primary evaluation originally relied on the MHJ dataset, which, though diverse in adversarial tactics, might not capture the full complexity of real-world multi-turn jailbreaks. Although we have extended our experiments to two additional datasets—ATTACK_600Ren et al. (2024) and CoSafeYu et al. (2024a)—these still represent curated benchmarks rather than open-ended, “in-the-wild” attacks. Future studies should expand testing to broader, unstructured scenarios (e.g., user-generated multi-turn dialogues on public platforms), ensuring that M2S generalizes across a more extensive range of adversarial strategies and topic domains.
  2. 2.
    Best-Case Performance Metric. In reporting M2S results, we often highlight the highest Attack Success Rate (ASR) across several formatting variants (e.g., Pythonize, Hyphenize, Numberize). Although this illustrates the upper bound of M2S effectiveness, it does not account for scenarios in which attackers cannot systematically select the optimal formatting strategy. Adversaries often adapt prompts iteratively, and real-time constraints may prevent them from exhaustively testing each M2S variant. Incorporating more realistic “on-the-fly” methods—such as automated tactic classifiers or adaptive prompt-generation agents—would provide a fuller picture of attainable, rather than purely best-case, performance.
  3. 3.
    Limited Automation of the Conversion Pipeline. Currently, we transform multi-turn dialogues into single-turn prompts (Hyphenize/Numberize/Pythonize) using rule-based templates. While these templates are deliberately simple, they require offline processing of the original multi-turn text. A fully automated system that:
    1. (a)
      identifies multi-turn adversarial clues in real time,
    2. (b)
      determines the most suitable single-turn “flattening” strategy,
    3. (c)
      re-injects the aggregated prompt into the LLM without human intervention
    would greatly advance both red-teaming and large-scale safety evaluations. Future work should develop end-to-end pipelines capable of online detection, dynamic formatting, and real-time performance measurement (e.g., with automated judges like StrongREJECT or JailbreakBench).
  4. 4.
    Potential Over-Reliance on a Single Automated Judge. Although StrongREJECT consistently aligns well with human annotations and offers a continuous harmfulness scale, there is still a risk of both false positives (classifying harmless text as harmful) and false negatives (missing subtly harmful content). The fixed threshold ($\mathrm{score}\geq 0.25$) was optimized via F1 scoring on specific labeled data, but different contexts might require alternative calibrations. Future research could triangulate multiple evaluation frameworks—including human raters and alternative automated judges—to confirm that M2S’s adversarial potency is robust across diverse safety metrics.
  5. 5.
    Model-Architecture Variability and Interpretability. Our findings indicate that certain M2S formats (e.g., “Pythonize”) excel with specific model families, whereas others (“Hyphenize”) can dominate on different architectures. This variability suggests that each LLM’s pretraining corpus and system prompts may parse structural cues differently. However, explaining why a given formatting style outperforms another on a given LLM remains difficult due to proprietary or black-box model internals. Deeper interpretability studies—potentially via sequence attributions or hidden-state analyses—could help clarify how enumerated versus code-like structures influence an LLM’s vulnerability to jailbreaks.
  6. 6.
    Real-World Complexity Beyond Token Counts. While our experiments show that M2S can drastically reduce token usage (and thus overhead) compared to multi-turn adversarial interactions, real-world deployments often involve additional system prompts, plugin calls, or reasoning traces that inflate total token counts. In a live chat system, for example, the token savings from M2S might be overshadowed by user-specific context or session-level state tokens. Future investigations should benchmark M2S in full, production-grade environments to gauge the practical cost savings and confirm that condensed single-turn prompts still bypass robust, real-time guardrails.
  7. 7.
    No Integrated Defense Proposals. Our study focuses primarily on revealing a new class of powerful single-turn threats rather than proposing comprehensive defenses. Although we highlight certain weaknesses—such as “contextual blindness” to enumerated or code-structured prompts—closing these gaps will likely require multi-layered approaches (e.g., improved token-level content filtering, heuristic detection of “flattened” prompts, or dynamic context-checking across simulated conversation states). Future lines of work should therefore develop and test defense strategies that operate effectively against M2S-style prompts, combining both static analysis (e.g., scanning for enumerations or code blocks) and runtime monitoring (e.g., verifying internal policy compliance per sub-request).
In summary, while our M2S framework underscores a critical vulnerability in current LLM safeguard designs and demonstrates how multi-turn adversarial dialogues can be “flattened” into potent single-turn prompts, a number of open questions remain. Addressing these—through broader datasets, more realistic best-case modeling, fully automated pipelines, multi-judge evaluation, interpretability research, real-world integration, and focused defensive strategies—will be essential for advancing the field of LLM safety and ensuring robust protection against ever-evolving adversarial prompt engineering.
Ethical Considerations
Our work aims to illuminate critical weaknesses in contemporary large-language-model (LLM) safety mechanisms by introducing methods that consolidate multi-turn adversarial tactics into single-turn prompts. Below we outline the key ethical considerations involved in designing and disseminating this research.
Intended use and potential misuse
The techniques we propose—Hyphenize, Numberize, and Pythonize—can reveal critical vulnerabilities in LLMs, but they could also be misapplied to generate harmful or disallowed content at scale. We release our methods and data to empower researchers and practitioners to identify and remediate these vulnerabilities, not to promote malicious exploitation. Any such use would contravene the spirit and intent of our work.
Data and content warnings
The datasets employed in this study (e.g. Multi-turn Human Jailbreak, Attack_600, CoSafe) inevitably contain adversarial prompts that request harmful or disallowed outputs from LLMs. We explicitly annotate these datasets as containing potentially sensitive, unsafe, or harmful text and strongly discourage their use outside carefully controlled safety evaluations. The data are released solely to facilitate the development and testing of robust defensive strategies.
Responsible release
We have taken care to limit the direct reproduction of explicitly harmful content in our examples and tables, focusing instead on high-level analyses of adversarial prompts and aggregate experiment results. Where possible, we redact or paraphrase content to minimise the risk of exposing sensitive or disallowed information. The consolidated single-turn prompts are made available only for reproducibility and to enable safety researchers to analyse emerging threats.
Broader impact on LLM safety
By demonstrating that single-turn, “flattened” adversarial prompts can sometimes surpass even carefully orchestrated multi-turn attacks, we aim to motivate the development of stronger guard-rail systems. In particular, our findings underscore the need to scrutinise entire blocks of input comprehensively rather than relying on turn-by-turn context checks alone. We hope that shedding light on these vulnerabilities will ultimately lead to more robust alignment and safer deployment of LLMs.
Overall, we adhere to the principle of coordinated disclosure by publishing our findings with clear disclaimers and minimal direct reproduction of harmful text. We hope this work will guide community-wide initiatives aimed at reinforcing LLM safety and mitigating the risks associated with adversarial attacks.
Appendix A Expanded Token Count Analysis
We have conducted a comprehensive assessment using OpenAI’s officially recommended tiktoken library (o200k_base) to better quantify token usage across different multi-turn jailbreak datasets. To ensure broader applicability beyond the MHJ dataset, we also ran additional experiments on two other multi-turn jailbreak datasets: ATTACK_600 and CoSafe. ATTACK_600 comprises especially lengthy multi-turn instructions crafted by advanced red-teamers, while CoSafe focuses on succinct scenario-driven adversarial queries. By including these diverse datasets, we capture a broader spectrum of conversation styles that challenge LLM guardrails from multiple angles.
Dataset Multi-turn Single-turn (M2S)
ATTACK_600 5331.32 830.61
MHJ 2732.24 1096.36
CoSafe 1689.28 469.10
Table 5: Average token counts measured with tiktoken (o200k_base) across three multi-turn jailbreak datasets.
These findings reinforce our assertion that the multi-turn format can accumulate significantly more tokens due to iterative message-passing between user and model. By contrast, a single-turn M2S prompt tends to be considerably shorter—yet can still retain (or even improve) the overall attack effectiveness, as demonstrated in Section 5.
Moreover, according to OpenAI’s official Conversation State guidelines, the total tokens consumed in an actual chat scenario can exceed these figures. This is because system instructions, internal reasoning tokens, or other role-based content may be appended automatically to maintain context across conversation turns. We hope these extended analyses offer a clearer perspective on how multi-turn interactions can inflate token usage, and why even shorter, single-turn prompts can remain highly adversarial.
Model Turn Method ASR (%) Average Score
GPT-4o-2024-11-20 Multi Original 57.3 0.53
Single Hyphenize (M2S) 49.3 (--8.0) 0.38 (--0.15)
Single Numberize (M2S) 49.3 (--8.0) 0.40 (--0.13)
Single Pythonize (M2S) 60.7 (+3.4) 0.56 (+0.03)
Single Ensemble (M2S) 67.0 (+9.7) 0.59 (+0.06)
Llama-3-70B-chat-hf Multi Original 46.0 0.37
Single Hyphenize (M2S) 38.7 (--7.3) 0.27 (--0.10)
Single Numberize (M2S) 39.3 (--6.7) 0.27 (--0.10)
Single Pythonize (M2S) 54.0 (+8.0) 0.42 (+0.05)
Single Ensemble (M2S) 57.0 (+11.0) 0.46 (+0.09)
Mistral-7B-Instruct-v0.3 Multi Original 63.7 0.46
Single Hyphenize (M2S) 75.7 (+12.0) 0.55 (+0.09)
Single Numberize (M2S) 74.3 (+10.6) 0.54 (+0.08)
Single Pythonize (M2S) 76.0 (+12.3) 0.56 (+0.10)
Single Ensemble (M2S) 86.0 (+22.3) 0.69 (+0.23)
GPT-4o-mini-2024-07-18 Multi Original 61.0 0.50
Single Hyphenize (M2S) 64.7 (+3.7) 0.48 (--0.02)
Single Numberize (M2S) 61.0 (+0.0) 0.45 (--0.05)
Single Pythonize (M2S) 72.0 (+11.0) 0.58 (+0.08)
Single Ensemble (M2S) 77.3 (+16.3) 0.64 (+0.14)
Table 6: M2S Performance on CoSafe. Attack-Success Rate (ASR) and Average StrongREJECT Score across models. Parentheses show the delta relative to the corresponding multi-turn baseline.
Appendix B M2S Performance on CoSafe
Below we report the M2S performance on the CoSafe dataset (see Table 6). The Hyphenize, Numberize, and Pythonize methods are applied to convert multi-turn adversarial queries into single-turn prompts. Overall, we observe that certain models see improvements in Attack Success Rate (ASR) with the Pythonize or ensemble approach, while others exhibit marginal decreases when using Hyphenize or Numberize alone. Notably, the “ensemble” metric (i.e., selecting the best result among the three M2S variants for each query) consistently outperforms the original multi-turn baseline.
Appendix C M2S Performance on ATTACK_600
Similarly, we show performance on the ATTACK_600 dataset (Table 7). Despite being relatively long and complex instructions, we again find that our M2S single-turn methods can retain or even boost overall ASR. The “Pythonize” variant often yields a substantial harmfulness increase, while “Hyphenize” and “Numberize” can still surpass the multi-turn baseline in some cases. Once more, the ensemble scenario demonstrates that consolidating the most effective single-turn result can provide a large improvement over the original multi-turn prompts.
Model Turn Method ASR (%) Average Score
GPT-4o-2024-11-20 Multi Original 71.5 0.61
Single Hyphenize (M2S) 69.0 (--2.5) 0.54 (--0.07)
Single Numberize (M2S) 67.8 (--3.7) 0.54 (--0.07)
Single Pythonize (M2S) 76.8 (+5.3) 0.69 (+0.08)
Single Ensemble (M2S) 91.0 (+19.5) 0.81 (+0.20)
Llama-3-70B-chat-hf Multi Original 63.8 0.42
Single Hyphenize (M2S) 71.0 (+7.2) 0.50 (+0.08)
Single Numberize (M2S) 63.8 (+0.0) 0.46 (+0.04)
Single Pythonize (M2S) 71.7 (+7.9) 0.46 (+0.04)
Single Ensemble (M2S) 92.2 (+28.4) 0.69 (+0.27)
Mistral-7B-Instruct-v0.3 Multi Original 73.2 0.42
Single Hyphenize (M2S) 80.2 (+7.0) 0.55 (+0.13)
Single Numberize (M2S) 80.5 (+7.3) 0.55 (+0.13)
Single Pythonize (M2S) 82.3 (+9.1) 0.54 (+0.12)
Single Ensemble (M2S) 96.5 (+23.3) 0.71 (+0.29)
GPT-4o-mini-2024-07-18 Multi Original 65.0 0.48
Single Hyphenize (M2S) 78.2 (+13.2) 0.56 (+0.08)
Single Numberize (M2S) 75.3 (+10.3) 0.53 (+0.05)
Single Pythonize (M2S) 74.2 (+9.2) 0.48 (+0.00)
Single Ensemble (M2S) 95.7 (+30.7) 0.72 (+0.24)
Table 7: M2S Performance on ATTACK_600. ASR and Average StrongREJECT Score for each model. Parentheses denote the change relative to the multi-turn baseline.
In summary, these additional datasets illustrate the generality of the M2S approach. Although multi-turn jailbreaks typically accumulate lengthy token contexts and can be extremely effective, flattening them into single-turn formats (while preserving the structural manipulations) can maintain or even enhance adversarial potency, often at a fraction of the token usage.
Appendix D Extended Tactic-Level Analysis
In addition to the primary evaluation metrics, we analyze adversarial tactics and how their StrongREJECT scores vary when transitioning from multi-turn to single-turn prompts. Specifically, we identify three types of shifts in harmfulness:
  • Score Increase Tactics: Jailbreak methods that become more harmful (from near 0 to $\geq 0.75$) when reformatted into a single-turn prompt.
  • Consistent High-Score Tactics: Methods that maintain a consistently high StrongREJECT score ($\geq 0.75$) across both multi-turn and single-turn scenarios.
  • Score Drop Tactics: Jailbreak methods that lose effectiveness (from $\geq 0.75$ to 0) when converted to single-turn formatting.
Tactic Score (↓) Appear
Irrelevant Distractor Instructions 1.73 12(39)
Suppressing Apologetic Behaviors 1.55 6(21)
Enforced Compliance to Harmful Command 1.27 23(82)
Legitimizing the Harmful Request with Positive, Affirmative Expressions 1.27 9(42)
Adding Distractor Instruction to Enforce Lexical/Syntactical Constraint 1.27 4(20)
Asking the Model in a Polite Tone 1.24 16(80)
Command to Ignore Previous Instructions 1.24 6(30)
Templated Output Format 1.15 43(226)
Potentially Rare Vanilla Harmful Request 1.13 27(143)
Elevating the Moral Grounding of a Harmful Request 1.10 21(115)
Enforced Rule-Breaking 1.07 10(55)
Irrelevant Distractor Components 1.03 30(167)
Providing Seed Examples 1.01 6(34)
Contextualizing the Task 1.01 81(463)
Leading Sentence Suffix 0.99 10(58)
Fabricate Moral Dilemma 0.88 13(77)
Downplaying the Request with More Nuanced Expressions 0.87 12(72)
Implied Harm 0.87 42(270)
Pretending 0.85 5(34)
Folding the Original Harmful Request into Another Nested Task 0.85 21(148)
Step-by-Step Instruction 0.75 16(122)
Assigning Model Personality 0.74 12(92)
Adding Distractor Instruction to Enforce Style Constraint 0.64 6(47)
Referring to Harmful Content by Pseudonym, Indirect Reference, or Coded Language 0.60 6(51)
Instructing the Model to Continue from the Refusal 0.58 2(18)
Surrogate Modality with Conversation 0.48 1(15)
Table 8: Score Increase Tactics: Jailbreak tactics that raise StrongREJECT scores from 0 to $\geq 0.75$ when moving from multi-turn to single-turn prompts.
As shown in Table 8, certain tactics—like Irrelevant Distractor Instructions and Suppressing Apologetic Behaviors—experience a pronounced jump in harmfulness when moved to a consolidated single-turn structure. This suggests that enumerating or embedding these cues in one prompt can amplify adversarial potency, possibly because the model treats the entire flattened request as a coherent directive, rather than fragmented instructions.
Tactic Score (↓) Appear
Assigning Model Personality 1.29 68(92)
Surrogate Modality with Conversation 1.23 11(15)
Referring to Harmful Content by Pseudonym, Indirect Reference, or Coded Language 1.23 34(51)
Adding Distractor Instruction to Enforce Style Constraint 1.21 31(47)
Folding the Original Harmful Request into Another Nested Task 1.19 97(148)
Pretending 1.18 22(34)
Legitimizing the Harmful Request with Positive, Affirmative Expressions 1.17 27(42)
Step-by-Step Instruction 1.14 76(122)
Templated Output Format 1.14 140(226)
Irrelevant Distractor Components 1.07 103(167)
Asking the Model in a Polite Tone 1.07 49(80)
Leading Sentence Suffix 1.05 35(58)
Contextualizing the Task 1.04 277(463)
Providing Seed Examples 1.04 20(34)
Implied Harm 1.03 155(270)
Elevating the Moral Grounding of a Harmful Request 1.01 64(115)
Instructing the Model to Continue from the Refusal 0.96 10(18)
Downplaying the Request with More Nuanced Expressions 0.90 40(72)
Fabricate Moral Dilemma 0.84 40(77)
Potentially Rare Vanilla Harmful Request 0.81 71(143)
Irrelevant Distractor Instructions 0.77 18(39)
Suppressing Apologetic Behaviors 0.64 9(21)
Enforced Compliance to Harmful Command 0.63 32(82)
Adding Distractor Instruction to Enforce Lexical/Syntactical Constraint 0.61 7(20)
Command to Ignore Previous Instructions 0.43 10(30)
Enforced Rule-Breaking 0.39 18(55)
Table 9: Consistent High-Score Tactics: Jailbreak tactics that keep StrongREJECT $\geq 0.75$ in both multi-turn and single-turn evaluations.
Meanwhile, Table 9 details those jailbreak patterns that remain consistently effective regardless of single-turn or multi-turn format. Techniques like Assigning Model Personality or Surrogate Modality with Conversation show consistently high StrongREJECT scores, indicating these strategies are robust to structural changes. Since such tactics already manipulate the model’s identity or rewrite context thoroughly, flattening them into a single turn likely does not dilute their adversarial content.
Tactic Score (↓) Appear
Instructing the Model to Continue from the Refusal 1.75 1(18)
Fabricate Moral Dilemma 1.71 4(77)
Assigning Model Personality 1.60 4(92)
Enforced Rule-Breaking 1.60 2(55)
Elevating the Moral Grounding of a Harmful Request 1.36 4(115)
Providing Seed Examples 1.32 1(34)
Potentially Rare Vanilla Harmful Request 1.30 4(143)
Implied Harm 1.28 7(270)
Irrelevant Distractor Instructions 1.26 1(39)
Irrelevant Distractor Components 1.24 4(167)
Legitimizing the Harmful Request with Positive, Affirmative Expressions 1.14 1(42)
Contextualizing the Task 1.14 11(463)
Adding Distractor Instruction to Enforce Style Constraint 1.07 1(47)
Folding the Original Harmful Request into Another Nested Task 1.06 3(148)
Leading Sentence Suffix 1.04 1(58)
Step-by-Step Instruction 0.92 2(122)
Downplaying the Request with More Nuanced Expressions 0.90 1(72)
Asking the Model in a Polite Tone 0.67 1(80)
Enforced Compliance to Harmful Command 0.65 1(82)
Templated Output Format 0.60 0(226)
Adding Distractor Instruction to Enforce Lexical/Syntactical Constraint 0.57 0(20)
Command to Ignore Previous Instructions 0.51 0(30)
Referring to Harmful Content by Pseudonym, Indirect Reference, or Coded Language 0.45 0(51)
Suppressing Apologetic Behaviors 0.35 0(21)
Pretending 0.33 0(34)
Surrogate Modality with Conversation 0.00 0(15)
Table 10: Score Drop Tactics: Jailbreak tactics whose StrongREJECT scores drop from $\geq 0.75$ to 00 when switching to single-turn prompts.
Lastly, Table 10 enumerates the jailbreak methods that appear to degrade in effectiveness once converted to single-turn. Strategies like Instructing the Model to Continue from the Refusal or Fabricate Moral Dilemma may rely heavily on the iterative, multi-turn structure—e.g., referencing an earlier refusal or building up user-manipulative context over multiple steps. Flattening these sequences into a single prompt may strip away the stepwise context or incremental justifications, reducing their harmfulness.
Overall, tactic-level analysis confirms that while many jailbreak cues can be preserved or even enhanced through single-turn formatting, a subset depends strongly on the dynamic, iterative nature of multi-turn interactions. Defenders should therefore examine both multi-turn and single-turn tactics, particularly those that show heightened susceptibility to flattening, when designing robust safety mechanisms.
Appendix E Real M2S Usage Example
Figure 2: An actual screenshot of a successful multi-turn jailbreak conversation, demonstrating how the user iteratively bypasses the model’s guardrails step by step. Each turn contributes additional manipulative cues that ultimately compel the model to produce harmful content.
Figure 3: An actual screenshot of the M2S (Multi-turn-to-Single-turn) usage, in which the entire multi-turn dialogue is consolidated into one prompt. Despite eliminating the iterative back-and-forth, the single-turn prompt can preserve — or even increase — the adversarial potency.
Appendix F Visualization Figures
Figure 4: This figure compares the Attack Success Rate (ASR) via bar charts for four models (GPT-4o, Llama-3-70B, Mistral-7B, GPT-4o-mini). Each model has five single-turn (M2S) techniques grouped by color (multi-turn, Hyphenize, Numberize, Pythonize, Ensemble). Overall, most models follow the trend Ensemble > individual M2S > multi-turn, with Pythonize alone often surpassing multi-turn significantly (notably on GPT-4o and Mistral-7B).
Figure 5: Using the same setup, we plot the “Perfect-ASR”—the fraction of responses scoring a perfect 1.0 on the harmfulness scale. The y-axis is 0–60%. Despite lower absolute values, the same pattern emerges: M2S consistently exceeds multi-turn, and Ensemble amplifies the effect, indicating that even the most harmful outputs can be achieved by single-turn reformatting.
Figure 6: Line plot of the average StrongREJECT (0–1) score, where higher values indicate more harmful content. Each line represents a single-turn technique; each marker represents a different model. Pythonize and Ensemble stand out above multi-turn, reflecting a tangible rise in the average level of harmfulness. Llama-3-70B is relatively lower with single M2S methods but returns to higher harmfulness with Ensemble.
Figure 7: A bar chart comparing the mean number of tokens per request on the MHJ dataset, contrasting the multi-turn approach (blue) with M2S (red). While multi-turn averages 2,732 tokens, M2S uses only 1,096 tokens, over a 60% reduction, yet still achieves higher or comparable ASR. This highlights that M2S can be both more efficient and more effective.
Figure 8: A bypass rate (“Safe” classification) comparison against the Llama-Guard-3-8B filter. While the single-turn M2S variants are similar or slightly lower than multi-turn individually, the Ensemble rises to 71%, surpassing even the multi-turn’s 66%. This indicates single-turn prompts can still mislead robust guardrails.
Figure 9: A row $\times$ column heatmap indicating “Which M2S transformation produced the highest harmfulness?” for each model. Darker cells indicate a larger share. We see model-specific preferences: GPT-4o-type and Mistral-7B strongly favor Pythonize; Llama-3-70B leans toward Hyphenize. The stark variation across models suggests that effective safety filters must adapt to individual model characteristics.
Figure 10: A bubble chart summarizing the top 10 adversarial tactics that show the largest increase in StrongREJECT scores after switching to single-turn. The x-axis measures the jump in harmfulness; the y-axis lists tactic names; bubble sizes represent frequency in the dataset. For instance, “Irrelevant Distractor Instructions” and “Suppressing Apologies” combine both frequent usage and major score increases, indicating that attackers may heavily exploit these methods.
Figure 11: A similar bubble chart for the 10 adversarial tactics that lose the most harmfulness score when consolidated into single-turn. The x-axis is the difference (multi-turn minus single-turn). Tactics such as “Continue from Refusal” and “Fabricate Moral Dilemma” strongly depend on multi-turn context, collapsing in potency when flattened into a single prompt.